# Data-driven decision-making under uncertainty with entropic risk measure

### v1

@@ 4.3 Numerical illustrations
Consider a project selection problem with three projects. Let $\xi\sim\text{GMM}\left({{\boldsymbol{\pi}}},{{\boldsymbol{\mu}}},{{\boldsymbol{\sigma}}}\right)$ with ${{\boldsymbol{\pi}}}\in\mathbb{R}^{5}$. The expected value of $\xi$ is $-18.6$ and standard deviation is $2.9$. Suppose the losses associated with the three projects are given by $0.4\xi$, $0.6\xi$ and $0.8\xi$, respectively. Let the risk aversion parameter be $\alpha=3$. The true entropic risk computed as: $\rho_{\mathbb{P}}(\xi)=\frac{1}{3}\log\left(\sum_{y=1}^{5}\pi_{y}\exp\left(3\mu_{y}+\frac{9}{2}\sigma_{y}^{2}\right)\right)$. To evaluate the estimates of entropic risk, we draw $1000$ instances with a sample size of $N=10000$ from the $\text{GMM}\left({{\boldsymbol{\pi}}},{{\boldsymbol{\mu}}},{{\boldsymbol{\sigma}}}\right)$.
SAA:
We compute the empirical entropic risk using the SAA for each project, resulting in the boxplots shown in Figure 4. The SAA method tends to underestimate the true risk because it fails to account adequately for tail events that have a low probability but significant impact on the entropic risk.
We can construct estimators for the entropic risk using leave-one-out cross-validation (LOOCV) and the Optimizer’s Information Criterion (OIC). Both approaches are based on an alternative representation of the entropic risk as an optimized certainty equivalent risk measure (Ben-Tal and Teboulle,, 1986):
where $h(t,\ell({{\boldsymbol{\xi}}}))=t+\phi(\ell({{\boldsymbol{\xi}}})-t),\phi(v)=\frac{1}{\alpha}\exp(\alpha v)-\frac{1}{\alpha}$. The optimal solution is given by $t^{*}=\rho_{{\mathbb{P}}}(\ell({{\boldsymbol{\xi}}}))$.
LOOCV:
The LOOCV estimator provides an alternative way to estimate entropic risk. For each sample ${{\boldsymbol{\xi}}}_{i}$, we solve the SAA problem without ${{\boldsymbol{\xi}}}_{i}$, obtaining
$\hat{t}_{{-i}}=\frac{1}{\alpha}\log\left(\frac{1}{N-1}\sum_{j\in[N]_{-i}}[\exp(\alpha\ell(\hat{{{\boldsymbol{\xi}}}}_{j}))]\right),$
where $[N]_{-i}=[N]\backslash\{i\}$. The LOOCV estimator is then given by
$\rho_{\text{loocv}}=\frac{1}{N}\sum_{i=1}^{N}h(\hat{t}_{-i},\ell({{\boldsymbol{\xi}}}))=\frac{1}{N}\sum_{i=1}^{N}\left(\hat{t}_{-i}+\frac{1}{\alpha}\exp(\alpha(\ell(\hat{{{\boldsymbol{\xi}}}}_{i})-\hat{t}_{-i}))-\frac{1}{\alpha}\right).$
Since $\hat{t}_{-i}$ is a feasible solution of the minimization problem in (8) for all $i\in[N]$, we have $\rho_{{\mathbb{P}}}(\ell({{\boldsymbol{\xi}}}))\leq{\mathbb{E}}_{{\mathbb{P}}}(h(\hat{t}_{-i},\ell({{\boldsymbol{\xi}}})))$ for each $i$. Thus, $\rho_{{\mathbb{P}}}(\ell({{\boldsymbol{\xi}}}))\leq{\mathbb{E}}_{{\mathbb{P}}}[\frac{1}{N}\sum_{i=1}^{N}h(\hat{t}_{-i},\ell({{\boldsymbol{\xi}}}))]={\mathbb{E}}_{{\mathbb{P}}}[\rho_{loocv}]$. We have show that $\rho_{\mbox{\scriptsize loocv}}$ is positively biased. Even though LOOCV estimators overestimate the true entropic risk, we can see in Figure 4 that the estimator has a skewed distribution where the entropic risk is below the true risk for more than $75\%$ times.
OIC:
Another approach to correct the first-order optimistic bias associated with solving SAA problem is given in Iyengar et al., (2023). They propose that SAA problem is solved with a decision rule $t(\theta)=\theta$ to obtain $\hat{\theta}=\rho_{\hat{{\mathbb{P}}}_{N}}(\ell({{\boldsymbol{\xi}}}))$. Then, the bias correction term is given by $\frac{1}{\alpha N^{2}}\sum_{i=1}^{N}\left(1-\exp(\alpha(\ell(\hat{{{\boldsymbol{\xi}}}}_{i})-\hat{\theta}))\right)^{2}$ (see Appendix B.2 for detailed calculations).
MoM.
The median-of-means (MoM) estimator partitions the loss scenarios into $B$ blocks, computes the mean within each block, and then takes the median of these means. This estimator is robust to outliers and is widely studied in the context of robust statistics (Lugosi and Mendelson,, 2019).
MLE:
We fit a GMM to the loss scenarios using the Expectation-Maximization algorithm, a popular method for maximizing the likelihood of the GMM parameters. However, the GMM fitted using MLE often underestimates the true entropic risk in the finite sample setting.
BS:
We apply the traditional bootstrapping approach (BS) by sampling 5000 observations with replacement from the empirical distribution, computing the empirical entropic risk, and repeating this procedure 1000 times. We report the statistics of the average entropic risk across these repetitions.
@@
We can see in Figure 4 that the estimators discussed above, including LOOCV, significantly underestimate the true entropic risk in most of the randomly generated instances for projects 2 and 3. On the other hand, our proposed estimators (BS-EVT and BS-Match) are able to mitigate the underestimation problem. Specifically, BS-EVT overestimates the entropic risk and BS-Match reduces the bias significantly compared to SAA.
@@ 5 Distributionally Robust Optimization
In data-driven optimization, the objective is to solve the following problem:
$\rho^{*}=\mathop{\rm min}_{{{\boldsymbol{z}}}\in\mathcal{Z}}\rho_{\mathbb{P}}(\ell({{\boldsymbol{z}}},{{\boldsymbol{\xi}}}))=\frac{1}{\alpha}\log\left(\mathbb{E}_{{\mathbb{P}}}[\exp(\alpha\ell({{\boldsymbol{z}}},{{\boldsymbol{\xi}}}))]\right),$ (9)
where $\mathcal{Z}\subseteq\mathbb{R}^{d}$ denotes the set of all feasible decisions, ${{\boldsymbol{z}}}\in\mathcal{Z}$ denotes a decision, ${{\boldsymbol{\xi}}}\in\mathbb{R}^{d}$ denotes the uncertain parameter, ${{\boldsymbol{\xi}}}\sim{\mathbb{P}}$, and the loss function $\ell$ depends on both ${{\boldsymbol{z}}}$ and ${{\boldsymbol{\xi}}}$. Let ${{\boldsymbol{z}}}^{*}$ denote the optimal decision obtained by solving (9).
For the optimal solution of problem (9) to be well defined, we make the following standard assumptions:
@@
We assume that
  1. 1.
    $\ell({{\boldsymbol{z}}},{{\boldsymbol{\xi}}})$ is convex in ${{\boldsymbol{z}}}$ for almost surely every ${{\boldsymbol{\xi}}}\in\Xi$.
  2. 2.
    $\ell({{\boldsymbol{z}}},{{\boldsymbol{\xi}}})$ is continuously differentiable in ${{\boldsymbol{z}}}\in\mathcal{Z}$ for almost surely every ${{\boldsymbol{\xi}}}\in\Xi$.
  3. 3.
@@
    For all ${{\boldsymbol{z}}}\in\mathcal{Z}$, the tails of $\ell({{\boldsymbol{z}}},{{\boldsymbol{\xi}}})$ are exponentially bounded:
    ${\mathbb{P}}(|\ell({{\boldsymbol{z}}},{{\boldsymbol{\xi}}})|>a)\leq G\exp(-a\alpha C)\quad\text{ for all }a\geq 0,$
    for some $G>0$ and $C>2$.
The true underlying distribution ${\mathbb{P}}$ is typically not known in practice, so one instead uses the samples $\mathcal{D}_{N}:=\{\hat{{{\boldsymbol{\xi}}}}_{1},\hat{{{\boldsymbol{\xi}}}}_{2},\cdots,\hat{{{\boldsymbol{\xi}}}}_{N}\}$ to solve the SAA of (9):
@@
$\rho_{\text{SAA}}=\mathop{\rm min}_{{{\boldsymbol{z}}}\in\mathcal{Z}}\rho_{\hat{{\mathbb{P}}}_{N}}(\ell({{\boldsymbol{z}}},{{\boldsymbol{\xi}}})):=\frac{1}{\alpha}\log\left(\mathbb{E}_{\hat{{\mathbb{P}}}_{N}}[\exp(\alpha\ell({{\boldsymbol{z}}},{{\boldsymbol{\xi}}}))]\right),$ (10)
where the empirical distribution, ^$\hat{{\mathbb{P}}}_{N}$, of the scenarios replaces the true distribution. Under fairly mild assumptions, the optimal risk for the SAA problem denoted by $\rho_{\text{SAA}}$ converges to the true risk $\rho^{*}$, see Lemma 2 in Appendix A. Thus, SAA approach is typically used to solve entropic risk minimization problems (Chen and Sim,, 2024); refer to Shapiro et al., (2009) for a comprehensive overview of SAA approach for stochastic programming.
In the limited data setting, the risk produced by solving problem (10) underestimates the true risk $\rho^{*}$ due to overfitting on the empirical distribution. Distributionally robust optimization (DRO) is one of the approaches to mitigate the optimistic bias of SAA by robustifying decisions against perturbations in the empirical distribution (Wiesemann et al.,, 2014; Delage and Ye,, 2010; Rahimian and Mehrotra,, 2022). It is assumed that nature perturbs the nominal distribution, e.g., empirical distribution, in a distributional ambiguity set $\mathcal{B}_{p}(\epsilon)$ so as to maximize the entropic risk of the decision maker, while decision maker aims to minimize the worst-case risk resulting in the following min-max problem:
@@
$\mathop{\rm min}_{{{\boldsymbol{z}}}\in\mathcal{Z}}\mathop{\rm sup}_{\mathbb{Q}\in\mathcal{B}_{p}(\epsilon)}\frac{1}{\alpha}\log\left(\mathbb{E}_{\mathbb{Q}}[\exp(\alpha\ell({{\boldsymbol{z}}},{{\boldsymbol{\xi}}}))]\right),$ (11)
where $\mathcal{B}_{p}(\epsilon)$ contains all distributions $\mathbb{Q}$ that are at a “distance” $\epsilon\geq 0$ away from the empirical distribution ^$\hat{{\mathbb{P}}}_{N}$. Here, $\epsilon$ is a parameter that controls the size of the ambiguity set and is chosen by the decision maker.
@@
In the literature, different ambiguity sets have been considered with the Kullback Leibler (KL)-divergence (Hu and Hong,, 2012) and Wasserstein ambiguity sets (Mohajerin Esfahani and Kuhn,, 2018) being the most commonly used (Rahimian and Mehrotra,, 2022). For KL-divergence-based ambiguity sets, the worst-case distribution is absolutely continuous with respect to the empirical distribution and does not contain distributions with support on points for which the empirical distribution is not supported. This makes the KL-divergence-based ambiguity set unsuitable for our problem with unbounded support. On the other hand, type-$p$ Wasserstein ambiguity set with $p<\infty$ results in unbounded worst-case loss when used in problem (11) (see Theorem 6 in Appendix B):
The type-$\infty$ Wasserstein distance is defined as follows:
@@
Let the ambiguity set $\mathcal{B}^{s}_{\infty}(\epsilon)$ be defined as:
$\mathcal{B}^{s}_{\infty}(\epsilon):=\left\{\mathbb{Q}\in\mathcal{M}(\Xi)\rvert\exists\,\,{{\boldsymbol{\xi}}}_{i}\in\Xi,\,\forall i\in[N],\|{{\boldsymbol{\xi}}}_{i}-\hat{{{\boldsymbol{\xi}}}}_{i}\|\leq\epsilon,\mathbb{Q}({{\boldsymbol{\xi}}})=\frac{1}{N}\sum_{i=1}^{N}\delta_{{{\boldsymbol{\xi}}}_{i}}({{\boldsymbol{\xi}}})\right\},$ (13)
where $\delta_{{\boldsymbol{\xi}}}$ is a Dirac measure at ${{\boldsymbol{\xi}}}$. From Bertsimas et al., (2023), problem (11) can be equivalently written as follows:
$\mathop{\rm min}_{{{\boldsymbol{z}}}\in\mathcal{Z}}\mathop{\rm sup}_{\mathbb{Q}\in\mathcal{B}^{s}_{\infty}(\epsilon)}\frac{1}{\alpha}\log\left(\mathbb{E}_{\mathbb{Q}}[\exp(\alpha\ell({{\boldsymbol{z}}},{{\boldsymbol{\xi}}}))]\right).$ (14)
Since $\log$ is a monotonic function, problem (14) is equivalent to:
$\mathop{\rm min}_{{{\boldsymbol{z}}}\in\mathcal{Z}}\frac{1}{\alpha}\log\left(\frac{1}{N}\sum_{i\in[N]}\mathop{\rm sup}_{{{\boldsymbol{\xi}}}_{i}:\|{{\boldsymbol{\xi}}}_{i}-\hat{{{\boldsymbol{\xi}}}}_{i}\|\leq\epsilon}\exp(\alpha\ell({{\boldsymbol{z}}},{{\boldsymbol{\xi}}}_{i}))\right).$ (15)
Thus, DRO problem in (11) with type-$\infty$ Wasserstein ambiguity set is equivalent to the robust optimization problem (15) with one uncertainty set associated with each scenario and nature can perturb each scenario $\hat{{{\boldsymbol{\xi}}}}_{i}$ in the ball $\|{{\boldsymbol{\xi}}}_{i}-\hat{{{\boldsymbol{\xi}}}}_{i}\|\leq\epsilon$. Thus, the worst-case loss is always bounded for finite radius $\epsilon$ of the uncertainty set.
Consider the loss function $\ell({{\boldsymbol{z}}},{{\boldsymbol{\xi}}}):={{\boldsymbol{z}}}^{\top}{{\boldsymbol{\xi}}}$. The function $\exp(\alpha{{\boldsymbol{z}}}^{\top}{{\boldsymbol{\xi}}})$ is increasing in ${{\boldsymbol{z}}}^{\top}{{\boldsymbol{\xi}}}$. Thus, its supremum over the set $\{{{\boldsymbol{\xi}}}_{i}:\|{{\boldsymbol{\xi}}}_{i}-\hat{{{\boldsymbol{\xi}}}}_{i}\|\leq\epsilon\}$ is attained at the boundary of this set. This leads to the result:
$\mathop{\rm sup}_{{{\boldsymbol{\xi}}}_{i}:\|{{\boldsymbol{\xi}}}_{i}-\hat{{{\boldsymbol{\xi}}}}_{i}\|\leq\epsilon}\exp(\alpha{{\boldsymbol{z}}}^{\top}{{\boldsymbol{\xi}}}_{i})=\exp\left(\alpha{{\boldsymbol{z}}}^{\top}\hat{{{\boldsymbol{\xi}}}}_{i}+\alpha\mathop{\rm sup}_{{{\boldsymbol{\xi}}}_{i}:\|{{\boldsymbol{\xi}}}_{i}\|\leq\epsilon}{{\boldsymbol{z}}}^{\top}{{\boldsymbol{\xi}}}_{i}\right)=\exp(\alpha{{\boldsymbol{z}}}^{\top}\hat{{{\boldsymbol{\xi}}}}_{i}+\epsilon\alpha\|{{\boldsymbol{z}}}\|_{*}),$ (16)
[…]
$\frac{1}{\alpha}\log\left(\frac{1}{N}\sum_{i=1}^{N}\exp(\alpha{{\boldsymbol{z}}}^{\top}\hat{{{\boldsymbol{\xi}}}}_{i}+\epsilon\alpha\|{{\boldsymbol{z}}}\|_{*})\right)=\frac{1}{\alpha}\log\left(\frac{1}{N}\sum_{i=1}^{N}\exp(\alpha{{\boldsymbol{z}}}^{\top}\hat{{{\boldsymbol{\xi}}}}_{i})\right)+\epsilon\left\lVert{{\boldsymbol{z}}}\right\rVert_{*}.$
So, with a linear loss function, problem (14) is equivalent to the regularized risk-averse SAA problem:
$\rho_{\text{DRO}}:=\mathop{\rm argmin}_{{{\boldsymbol{z}}}\in\mathcal{Z}}\frac{1}{\alpha}\log\left(\frac{1}{N}\sum_{i=1}^{N}\exp(\alpha{{\boldsymbol{z}}}^{\top}\hat{{{\boldsymbol{\xi}}}}_{i})\right)+\epsilon\left\lVert{{\boldsymbol{z}}}\right\rVert_{*}.$ (17)
For piecewise concave loss functions, Theorem 2 gives an equivalent reformulation of the DRO problem as the finite dimensional convex optimization problem using Fenchel duality.
@@
$\mathop{\rm minimize}_{{{\boldsymbol{t}}},\,{{\boldsymbol{z}}}\in\mathcal{Z},\,\{\boldsymbol{\varphi_{ij}}\}_{i\in[N],\,j\in[m]}}$ $\quad\frac{1}{\alpha}\log\left(\frac{1}{N}\sum_{i=1}^{N}\exp(\alpha t_{i})\right)$ (18a)
subject ${\rm subject\mbox{ }\rm to}$ $\quad\hat{{{\boldsymbol{\xi}}}}_{i}^{\top}{{\boldsymbol{\varphi}}}_{ij}+\epsilon\|{{\boldsymbol{\varphi}}}_{ij}\|_{*}-\ell_{j*}({{\boldsymbol{z}}},{{\boldsymbol{\varphi}}}_{ij})\leq t_{i}\,,\;\forall i\in[N],\,j\in[m],$ (18b)

### v3

@@ 4.3 Comparison with benchmarks from the literature
Consider a project selection problem with three projects. Let $\xi\sim\text{GMM}\left({{\boldsymbol{\pi}}},{{\boldsymbol{\mu}}},{{\boldsymbol{\sigma}}}\right)$ with $Y=5$ (see Appendix B.5 for parameter values of the GMM with randomly generated weights $\boldsymbol{\pi}$). Suppose the losses associated with the three projects are given by $0.4\xi$, $0.6\xi$ and $0.8\xi$, respectively. Let the risk aversion parameter be $\alpha=3$. The true entropic risk can be analytically expressed as $\rho_{\mathbb{P}}(\xi)=\frac{1}{3}\log\left(\sum_{y=1}^{Y}\pi_{y}\exp\left(3\mu_{y}+\frac{9}{2}\sigma_{y}^{2}\right)\right)$. To evaluate each estimator, we draw $1000$ instances with a sample size of $N=10000$ from the $\text{GMM}\left({{\boldsymbol{\pi}}},{{\boldsymbol{\mu}}},{{\boldsymbol{\sigma}}}\right)$.
@@
Figure 4 shows that the estimators discussed above significantly underestimate the true entropic risk for projects 2 and 3. It is interesting to note that although the LOOCV estimator overestimates the true entropic risk, its distribution across the $1000$ instances is skewed, with the entropic risk falling below the true value in over 75% of instances. In contrast, our proposed estimators effectively address the underestimation issue. Specifically, BS-EVT overestimates the entropic risk and BS-Match reduces the bias of SAA estimator significantly.
@@
In the next section, we will show that the proposed procedures for mitigating estimation bias can also be applied to mitigate the optimistic bias when solving entropic risk minimization problems. As discussed earlier, optimistic bias occurs due to lack of data, in which case regularization type techniques (such as distributionally robust optimization) are employed to correct the bias. By providing better estimates of the validation risk, we can more accurately calibrate the hyperparameters compared to traditional CV methods.
@@ 5 Distributionally Robust Optimization
Entropic risk minimization considers the following problem:
$\rho^{*}=\mathop{\rm min}_{{{\boldsymbol{z}}}\in\mathcal{Z}}\rho_{\mathbb{P}}(\ell({{\boldsymbol{z}}},{{\boldsymbol{\xi}}})):=\frac{1}{\alpha}\log\left(\mathbb{E}_{{\mathbb{P}}}[\exp(\alpha\ell({{\boldsymbol{z}}},{{\boldsymbol{\xi}}}))]\right),$ (7)
where $\mathcal{Z}\subseteq\mathbb{R}^{d}$ denotes the set of all feasible decisions, ${{\boldsymbol{\xi}}}\in\mathbb{R}^{d}$ denotes the uncertain vector following the probability distribution ${\mathbb{P}}$, and $\ell(\boldsymbol{z},\boldsymbol{\xi})$ denotes the loss function. For the optimal solution $\boldsymbol{z}^{*}$ of problem (7) to be well defined, we make the following standard assumptions:
@@
We assume that:
  1. (A.1)
    $\mathcal{Z}$ is a compact and convex set.
  2. (A.2)
    $\ell({{\boldsymbol{z}}},{{\boldsymbol{\xi}}})$ is convex in ${{\boldsymbol{z}}}$ for almost every ${{\boldsymbol{\xi}}}\in\Xi$.
  3. (A.3)
    $\ell({{\boldsymbol{z}}},{{\boldsymbol{\xi}}})$ is $L$-Lipschitz continuous in ${{\boldsymbol{\xi}}}$ for all ${{\boldsymbol{z}}}\in\mathcal{Z}$.
  4. (A.4)
    $\ell({{\boldsymbol{z}}},{{\boldsymbol{\xi}}})$ is $L({{\boldsymbol{\xi}}})$-Lipschitz continuous in ${{\boldsymbol{z}}}$ for all ${{\boldsymbol{\xi}}}\in\Xi$ with ${\mathbb{E}}_{\mathbb{P}}[L({{\boldsymbol{\xi}}})^{q}]<\infty$ for all $q\geq 1$.
  5. (A.5)
@@
    $|\ell({{\boldsymbol{z}}},{{\boldsymbol{\xi}}})|\leq\bar{L}({{\boldsymbol{\xi}}})$ for all ${{\boldsymbol{z}}}\in\mathcal{Z}$ almost surely, with the tail of $\bar{L}({{\boldsymbol{\xi}}})$ exponentially bounded:
    ${\mathbb{P}}\big{(}\bar{L}({{\boldsymbol{\xi}}})>a\big{)}\leq G\exp(-a\alpha C)\quad\text{for all }a\geq 0,$
    for some constants $G>0$ and $C>2$.
As in Assumption 2, the last assumption ensures that the mean and variance of the loss $\exp(\alpha\ell({{\boldsymbol{z}}},{{\boldsymbol{\xi}}}))$ are finite for each ${{\boldsymbol{z}}}\in\mathcal{Z}$. We make the technical assumptions (A.4) and (A.5) to ensure the convergence of SAA solution to the true risk (see Proposition 3).
As the true underlying distribution ${\mathbb{P}}$ is typically unknown, it is common practice to replace it with the empirical distribution ^$\hat{\mathbb{P}}_{N}$, solving the corresponding SAA problem:
@@
$\rho_{\text{{SAA}}}=\mathop{\rm min}_{{{\boldsymbol{z}}}\in\mathcal{Z}}\rho_{\hat{{\mathbb{P}}}_{N}}(\ell({{\boldsymbol{z}}},{{\boldsymbol{\xi}}})):=\frac{1}{\alpha}\log\left(\mathbb{E}_{\hat{{\mathbb{P}}}_{N}}[\exp(\alpha\ell({{\boldsymbol{z}}},{{\boldsymbol{\xi}}}))]\right).$ (8)
Under Assumption 3, it follows that $\rho_{\text{{SAA}}}\rightarrow\rho^{*}$ as $N$ grows, as shown in Proposition 3.
Proposition 3
Suppose that Assumption 3 holds. Then, for any $\gamma>0$, there exists a constant $A>0$ for which,
@@
${\mathbb{P}}$ $\left(\left|\rho_{\text{{SAA}}}-\rho^{*}\right|\geq\frac{A}{\sqrt{N\gamma}\alpha\exp(\alpha\rho^{*})}\right)\leq\gamma,$ (9)
as long as $N$ is sufficiently large. Consequently, $\rho_{\text{{SAA}}}\rightarrow\rho^{*}$ in probability.
To prove Proposition 3, we show that $\exp(\alpha\ell({{\boldsymbol{\xi}}},{{\boldsymbol{z}}}))$ is $\kappa({{\boldsymbol{\xi}}})$-Lipschitz continuous in ${{\boldsymbol{z}}}$ for all ${{\boldsymbol{\xi}}}\in\Xi$ with $\kappa({{\boldsymbol{\xi}}})=\alpha L({{\boldsymbol{\xi}}})\exp(\alpha\bar{L}({{\boldsymbol{\xi}}}))$. Then, we apply the uniform convergence results for heavy tailed distributions in Jiang et al., (2020, Theorem 3.2) to show the uniform convergence of empirical utility to true utility for all ${{\boldsymbol{z}}}\in\mathcal{Z}$. This theorem only requires that the second moment of $\kappa({{\boldsymbol{\xi}}})$ is finite instead of the usual light-tailed assumptions that don’t hold for $\kappa({{\boldsymbol{\xi}}})$. Subsequently, we use the properties of logarithm function in the neighborhood of zero to show the uniform convergence of empirical risk to optimal risk. This convergence result underpins the prevalent use of the SAA approach for entropic risk minimization problems (Chen et al., 2024a, ; Chen et al., 2024b, ). For a detailed examination of the SAA methodology within stochastic programming, see Shapiro et al., (2009).
In the limited data setting, the risk produced by solving problem (8) underestimates the true risk $\rho^{*}$ due to overfitting on the empirical distribution. Distributionally robust optimization (DRO) is one of the approaches to mitigate the optimistic bias of SAA by robustifying decisions against perturbations in the empirical distribution (Delage and Ye,, 2010; Wiesemann et al.,, 2014; Mohajerin Esfahani and Kuhn,, 2018; Rahimian and Mehrotra,, 2022). It is assumed that nature perturbs the empirical distribution within a distributional ambiguity set $\mathcal{B}(\epsilon)$ containing all distributions $\mathbb{Q}$ that are at a “distance” $\epsilon\geq 0$ away from the empirical distribution ^$\hat{{\mathbb{P}}}_{N}$, so as to maximize the entropic risk of the decision maker, while decision maker aims to minimize the worst-case risk resulting in the following min-max problem:
$\rho_{\texttt{DRO}}:=\mathop{\rm min}_{{{\boldsymbol{z}}}\in\mathcal{Z}}\mathop{\rm sup}_{\mathbb{Q}\in\mathcal{B}(\epsilon)}\frac{1}{\alpha}\log\left(\mathbb{E}_{\mathbb{Q}}[\exp(\alpha\ell({{\boldsymbol{z}}},{{\boldsymbol{\xi}}}))]\right).$ (10)
The size of the ambiguity set, $\epsilon$, is chosen by the decision maker; however, as we will discuss later, in practice it is treated as a hyperparameter tuned through CV to optimize the performance of the optimization model on unseen data.
@@
In the literature, different ambiguity sets have been considered with the Kullback Leibler (KL)-divergence (Hu and Hong,, 2012) and Wasserstein ambiguity sets (Mohajerin Esfahani and Kuhn,, 2018) being the most commonly used (Rahimian and Mehrotra,, 2022). For KL-divergence-based ambiguity sets, the standard formulation (Hu and Hong,, 2012) restricts the worst-case distribution to be absolutely continuous with respect to the empirical distribution, limiting its support to the same points as the empirical distribution. This poses a problem because it prevents the representation of worst-case scenarios that typically occur in the tails of the loss distribution. An alternative formulation does allow worst-case distributions with support beyond the empirical distribution, enabling a richer ambiguity set (Chan et al.,, 2024). However, with an unbounded loss function, this flexibility allows nature to exploit the tail, resulting in infinite loss for the decision maker. Consequently, KL-divergence-based ambiguity sets are ill-suited to our problem, which involves unbounded support and heavy-tailed losses. This also holds for type-$p$ Wasserstein ambiguity set with $p<\infty$ due to the following result.
Proposition 4
The $p$-Wasserstein DRO with entropic risk measure results in unbounded loss if $p<\infty$.
Next, we show that type-$\infty$ Wasserstein ambiguity set is a suitable choice for problem (10). The type-$\infty$ Wasserstein distance is defined as:
@@
Bertsimas et al., (2023) have shown that problem (10) can be equivalently written as:
$\mathop{\rm min}_{{{\boldsymbol{z}}}\in\mathcal{Z}}\mathop{\rm sup}_{\mathbb{Q}\in\tilde{\mathcal{B}}_{\infty}(\epsilon)}\frac{1}{\alpha}\log\left(\mathbb{E}_{\mathbb{Q}}[\exp(\alpha\ell({{\boldsymbol{z}}},{{\boldsymbol{\xi}}}))]\right)=\mathop{\rm min}_{{{\boldsymbol{z}}}\in\mathcal{Z}}\frac{1}{\alpha}\log\left(\frac{1}{N}\sum_{i\in[N]}\mathop{\rm sup}_{{{\boldsymbol{\xi}}}:\|{{\boldsymbol{\xi}}}-\hat{{{\boldsymbol{\xi}}}}_{i}\|\leq\epsilon}\exp(\alpha\ell({{\boldsymbol{z}}},{{\boldsymbol{\xi}}}))\right),$ (12)
where the ambiguity set ~$\tilde{\mathcal{B}}_{\infty}(\epsilon)$ is defined as:
~$\tilde{\mathcal{B}}_{\infty}(\epsilon):=\left\{\mathbb{Q}\in\mathcal{M}(\Xi)\rvert\exists\,\,{{\boldsymbol{\xi}}}_{i}\in\Xi,\|{{\boldsymbol{\xi}}}_{i}-\hat{{{\boldsymbol{\xi}}}}_{i}\|\leq\epsilon,\,\forall i\in[N],\,\mathbb{Q}({{\boldsymbol{\xi}}})=\frac{1}{N}\sum_{i=1}^{N}\delta_{{{\boldsymbol{\xi}}}_{i}}({{\boldsymbol{\xi}}})\right\}.$
The equality in (12) follows since the logarithm function is monotone. From the formulation of problem (12), we can clearly see the behavior of the adversarial uncertainty, which allows each scenario $\hat{{{\boldsymbol{\xi}}}}_{i}$ to be perturbed within a norm ball of radius $\epsilon$. Thus, by construction, the worst-case loss is always bounded for finite radius $\epsilon$ of the uncertainty set.
For piecewise concave loss functions, the following theorem gives an equivalent reformulation of the DRO problem as the finite dimensional convex optimization problem using Fenchel duality (Ben-Tal et al.,, 2015).
@@
$\begin{array}[]{lll}\mathop{\rm min}&\frac{1}{\alpha}\log\left(\frac{1}{N}\sum_{i=1}^{N}\exp(\alpha t_{i})\right)\\ \text{s.t.}&\boldsymbol{t}\in\mathbb{R}^{N},\,\boldsymbol{z}\in\mathcal{Z},\,\boldsymbol{\varphi}_{ij}\in\mathbb{R}^{d}&\forall i\in[N],\,j\in[m]\\ &{{\boldsymbol{\varphi}}}_{ij}^{\top}\hat{{{\boldsymbol{\xi}}}}_{i}-\ell_{j*}({{\boldsymbol{z}}},{{\boldsymbol{\varphi}}}_{ij})+\epsilon\|{{\boldsymbol{\varphi}}}_{ij}\|_{*}\leq t_{i}&\forall i\in[N],\,j\in[m],\end{array}$ (13)